# TRec: Egocentric Action Recognition using 2D Point Tracks

### v2

@@ Abstract
We present a novel approach for egocentric action recognition that leverages 2D point tracks as an additional motion cue. While most existing methods rely on RGB appearance, human pose estimation, or their combination, our work demonstrates that tracking randomly sampled image points across video frames can substantially improve recognition accuracy. Unlike prior approaches, we do not detect hands, objects, or interaction regions. Instead, we employ CoTracker to follow a set of randomly initialized points through each video and use the resulting trajectories, together with the corresponding image frames, as input to a Transformer-based recognition model. Surprisingly, our method achieves notable gains even when only the initial frame and its associated point tracks are provided, without incorporating the full video sequence. Experimental results confirm that integrating 2D point tracks consistently enhances performance compared to the same model trained without motion information, highlighting their potential as a lightweight yet effective representation for egocentric action understanding.
@@ 1 Introduction
The task of action recognition involves identifying and classifying human actions from video data. In the egocentric setting, videos are captured from a first-person perspective, typically using head-mounted cameras. Unlike exocentric (third-person) videos, which offer a stable, external view of actors performing generic activities such as sports or daily tasks, egocentric videos focus on fine-grained, domain-specific actions centered around hand-object interactions and the wearer’s visual attention. This perspective introduces unique challenges, including strong camera motion, frequent occlusions, and limited field of view, all of which complicate the modeling of both actor and object dynamics.
The prominent camera motion paired with the exceptional close and fast-moving background covering an extensive amount of the image pose an intricate challenge of modelling both actor and object dynamics. The interplay of intense camera motion combined with the rapidly moving and closely captured background dominating the visual field make it particularly challenging to model both actor and object dynamics. Many existing approaches rely on RGB-based features. While such methods have achieved strong results, they often depend on computationally expensive or error-prone semantic detection pipelines for hands, objects, or skeletons.
In this work, we investigate whether simple, semantically agnostic motion cues, specifically random 2D point tracks, can provide comparable or even superior information for recognizing actions in egocentric videos.
By tracking multiple points on hands, objects, and background regions, our method captures diverse motion signals, including head movement and object deformation. The difficulty of distinguishing object categories is reduced, since our method directly tracks the motion of the object in the scene. We generate these tracks using CoTracker. The classification pipeline uses the ResNet18 as backbone and a Transformer based module on top for classification. Surprisingly, our experiments show that a Transformer model trained with such point tracks, even when only the initial frame is used, can recognize actions more accurately than the RGB-only baseline using the whole video.
@@
  • We propose a simple and efficient method for egocentric action recognition that integrates 2D point tracks with image features.
@@
  • We show consistent performance improvements over the RGB-only baseline on an egocentric benchmark.
@@ 2 Related Work
Trajectory-based methods model motion by tracking key points or regions across consecutive frames [18, 12, 20]. They have been explored in exocentric (third-person) action recognition, where stable camera viewpoints simplify trajectory estimation. To the best of our knowledge, no prior work has investigated the use of 2D point tracks for egocentric action recognition without filtering background points, where strong camera motion and frequent occlusions make trajectory-based modeling more challenging. Additionally, in contrast to these methods we utilize a relatively sparse set of points to track within the image. Wang et al. [18] introduced Improved Dense Trajectories, which compensate for camera motion using RANSAC-based estimation and employ a human detector to distinguish foreground motion from background motion. This approach established a strong pre-deep-learning baseline for video action recognition. Nguyen et al. [12] proposed Denser Trajectories of Anchor Points for Action Recognition, which anchor trajectories to more stable points to better capture fine-grained motion variations and increase robustness to camera movement.
@@
Hand-object interaction modeling can improve egocentric action understanding. Several methods utilize some kind of hand-object detector and integrate this information into their model [3, 14, 13, 11]. Chatterjee et al. [3] expanded the vocabulary of egocentric actions by linking detected hands and objects with fine-grained action categories. Shiota et al. [14] explicitly modeled hand-object contact and object state transitions, while Peirone et al. [13] proposed a zone-aware recognition frameworks which learns environmental affordances to extract environment and action information seperately to train an action recognition model. Materzyńska et al. [11] proposed Something-Else, a compositional model that builds spatial-temporal interaction graphs between hands and objects, achieving strong performance on the Something-Something-v2 dataset. Although effective, these methods depend heavily on accurate hand and object detection, which can fail under occlusions or strong motion.
@@
RGB-based approaches [15, 1, 9, 5, 16, 19] rely solely on raw visual appearance for action recognition, without incorporating explicit motion or geometric cues. Many of these high-performing models are trained on large-scale datasets, often leveraging extensive pretraining and carefully designed architectures. In contrast, we deliberately adopt a simple model design to isolate and clearly demonstrate the effectiveness of incorporating 2D point tracks. Unlike RGB-based features that primarily encode texture and color statistics, point trajectories provide explicit geometric and temporal information, capturing object and camera motion directly from the image sequence. Girdhar et al. [5] propose the Anticipative Video Transformer, which models temporal dependencies in egocentric videos to predict upcoming actions. Their model implicitly attends to image regions containing hands and objects, and its performance improves with increased temporal context. However, the paper does not clarify whether motion is inferred implicitly through appearance changes between frames. In contrast, our approach introduces motion information explicitly by integrating 2D point tracks, enabling the model to reason directly about the displacement and dynamics of scene elements rather than relying on learned appearance differences. This explicit representation of motion complements the visual features and enhances interpretability in modeling fine-grained egocentric actions. Tong et al. [16] employ masked autoencoder pretraining to improve video representation learning, achieving strong performance on egocentric and other video benchmarks. Wang et al. [19] further extend pretrained video backbone models through an interactive prototype learning framework, enhancing active-object representations by leveraging actor motion cues within RGB inputs.
@@
Point tracking aims to estimate the trajectories of 2D points across video frames, providing fine-grained motion information that supports higher-level tasks such as action recognition and object interaction analysis. Recent advances in this area have focused on improving both tracking accuracy and computational efficiency. CoTracker [8] introduced a transformer-based architecture capable of jointly tracking multiple points across frames. Subsequent versions have further improved robustness, accuracy, and inference speed, establishing CoTracker as a state-of-the-art solution for general-purpose point tracking. In our work, we adopt CoTracker due to the public availability of pretrained weights, which allows for efficient and reliable integration of high-quality point trajectories into our framework. EgoPoints [4] proposed a dedicated benchmark for dense point tracking in egocentric videos. This work specifically addresses the challenges of egocentric settings, such as strong camera motion and frequent occlusions, and provides a comprehensive evaluation protocol for tracking performance. In addition, the authors improve upon CoTracker by adapting it to egocentric motion patterns and by providing a large-scale dataset for benchmarking.
@@ 3 Action Recognition using Point Tracks
We present TRec, a transformer-based model for egocentric action recognition that jointly leverages 2D point tracks and image representations. TRec tracks both background and foreground points throughout a video to capture rich motion cues without relying on explicit object or hand detections. The model employs a lightweight image encoder to extract spatial features from individual frames and a transformer architecture [17] to integrate temporal dependencies across frames. A set of randomly sampled 2D points is initialized within each video, covering both foreground and background regions, and their trajectories are tracked over time. The transformer receives as input the encoded image features together with these 2D point tracks, enabling it to reason jointly about appearance and motion information. An additional multi-headed attention layer is used to aggregate the transformer outputs, and the resulting representation is passed through an MLP-based action recognition head to predict the action depicted in the video. We evaluate TRec on the Something-Something-v2 [6] dataset.
@@ 3.1 Dataset Curation and Preprocessing
Something-Something-v2. The Something-Something-v2 dataset [6] contains short video clips depicting human-object interactions recorded in a wide variety of everyday environments. It comprises 174 action templates describing generic activities of someone doing ”something”, such as pushing, pulling, or placing objects. It offers greater diversity in both visual context and action types, as well as a substantially large number of annotated samples, making it a valuable benchmark for training data-driven models. Because the camera remains largely stationary across most recordings, the dataset provides a stable visual setting that facilitates the analysis of object and hand dynamics without the confounding effects of egocentric motion. We use the dataset in its original form without any additional preprocessing or modifications.
Point Track Generation. We curate 2D point trajectories using CoTracker3 [7]. For each action sequence, we randomly sample 900 points covering the entire image, including both foreground and background regions. Video frames are sampled at 30 fps to obtain smooth trajectories. The resulting point coordinates are normalized following the procedure described in [2]. For each video we track a set of 900 points $p^{i}=\{(x^{i}_{t},y^{i}_{t})\in\mathbb{R}^{2}|t=0,...T\}$. Here, $i\in[0,P]$ denotes the $i^{\text{th}}$ point with $P$ being the total number of points while $t\in[0,T]$ refrs to the frame index, with $T$ denoting the number of frames in the video. We refer to a point at time step $t$ as $p_{t}^{i}\in\mathbb{R}^{2}$. The set of all point tracks within a video is denoted as $\mathbf{P}$.
(a) Hitting steel glass with screw driver
(b) Lifting a lighter and dropping it
(c) moving scissors and screwdriver closer
Figure 1: These images show the tracked trajectory of 2D points in videos of the Something-Something dataset.
@@ 3.2 Network Architecture
3.3 Action Recognition
The transformer model receives as input the image features extracted by the encoder together with the corresponding 2D point tracks. This joint representation allows the model to integrate appearance and motion information in a unified latent space. Each frame contributes both spatial features from the image encoder and geometric motion cues from the tracked points. By processing these modalities jointly, the transformer learns to associate appearance changes with underlying motion dynamics, enabling robust recognition of fine-grained actions in egocentric settings.
@@ 3.3 Action Recognition
Finally, the pooled class token is passed through a lightweight multi-layer perceptron (MLP) classification head to produce the final action prediction. During training, the model is optimized using a standard cross-entropy loss over the action classes defined by each dataset. This simple yet effective design allows TRec to capture meaningful motion representations directly from tracked points and image features, achieving competitive performance while maintaining low computational complexity.
@@ 4.1 Track vs no Track
Table 1: Track vs No-Track
@@ 4.3 Background Motion as Context for Action Recognition
Table 3: Model evaluation with Kernel Density Function
@@ 5 Discussion
The experimental results demonstrate that incorporating 2D point tracks provides a clear advantage for egocentric action recognition compared to purely RGB-based methods. By explicitly encoding motion cues through tracked points, our model captures spatial-temporal dynamics that RGB alone cannot represent. This finding supports our initial hypothesis that motion trajectories offer a lightweight and interpretable alternative to computationally demanding approaches such as optical flow or hand–object detection.

### v3

# TRec: Learning Hand-Object Interactions through 2D Point Track Motion

@@ Abstract
We present a novel approach for hand-object action recognition that leverages 2D point tracks as an additional motion cue. While most existing methods rely on RGB appearance, human pose estimation, or their combination, our work demonstrates that tracking randomly sampled image points across video frames can substantially improve recognition accuracy. Unlike prior approaches, we do not detect hands, objects, or interaction regions. Instead, we employ CoTracker to follow a set of randomly initialized points through each video and use the resulting trajectories, together with the corresponding image frames, as input to a Transformer-based recognition model. Surprisingly, our method achieves notable gains even when only the initial frame and the point tracks are provided, without incorporating the full video sequence. Experimental results confirm that integrating 2D point tracks consistently enhances performance compared to the same model trained without motion information, highlighting their potential as a lightweight yet effective representation for hand-object action understanding.
Keywords: action recognition, computer vision, motion-based recognition.
@@ 1 Introduction
The task of action recognition involves identifying and classifying human actions from video data. While many prior works focus on generic third-person datasets that capture large-scale body movements, our study targets hand-object interaction scenarios, where meaningful actions are characterized by fine-grained manipulations rather than full-body motion. Such datasets can include both egocentric and third-person viewpoints, providing diverse perspectives on close-range interactions between hands and objects. This perspective introduces unique challenges, including strong camera motion, frequent occlusions, and limited field of view, all of which complicate the modeling of both actor and object dynamics.
The prominent camera motion paired with the exceptional close and fast-moving background covering an extensive amount of the image pose an intricate challenge of modeling both actor and object dynamics. The interplay of camera motion combined with the rapidly moving and closely captured background dominating the visual field make it particularly challenging to model both actor and object dynamics. Many existing approaches rely on RGB-based features. While such methods have achieved strong results, they often depend on computationally expensive or error-prone semantic detection pipelines for hands, objects, or skeletons.
In this work, we investigate whether simple, semantically agnostic motion cues, specifically random 2D point tracks, can provide comparable or even superior information for recognizing hand-object interactions.
By tracking multiple points on hands, objects, and background regions, our method captures diverse motion signals, including camera movement and object deformation. The difficulty of distinguishing object categories is reduced, since our method directly tracks the motion of the object in the scene. We generate these tracks using CoTracker. The classification pipeline uses the ResNet18 as backbone and a Transformer based module on top for classification. Surprisingly, our experiments show that a Transformer model trained with such point tracks, even when only the initial frame is used, can recognize actions more accurately than the RGB-only baseline using the whole video.
@@
  • We propose a simple and efficient method for hand-object action recognition that integrates 2D point tracks with image features.
@@
  • We show consistent performance improvements over the RGB-only baseline on an hand-object interaction benchmark.
@@ 2 Related Work
Trajectory-based methods model motion by tracking key points or regions across consecutive frames [18, 12, 20]. They have been explored in exocentric (third-person) action recognition, where stable camera viewpoints simplify trajectory estimation. To the best of our knowledge, no prior work has investigated the use of 2D point tracks for hand-object action recognition without filtering background points, where strong camera motion and frequent occlusions make trajectory-based modeling more challenging. Additionally, in contrast to these methods we utilize a relatively sparse set of points to track within the image compared to methods using optical flow approaches. Wang et al. [18] introduced Improved Dense Trajectories, which compensate for camera motion using RANSAC-based estimation and employ a human detector to distinguish foreground motion from background motion. This approach established a strong pre-deep-learning baseline for video action recognition. Nguyen et al. [12] proposed Denser Trajectories of Anchor Points for Action Recognition, which anchor trajectories to more stable points to better capture fine-grained motion variations and increase robustness to camera movement.
@@
Hand-object detectors can improve hand-object interaction recognition. Several methods utilize some kind of hand-object detector and integrate this information into their model [3, 14, 13, 11]. Chatterjee et al. [3] expanded the vocabulary of egocentric actions by linking detected hands and objects with fine-grained action categories. Shiota et al. [14] explicitly modeled hand-object contact and object state transitions, while Peirone et al. [13] proposed a zone-aware recognition frameworks which learns environmental affordances to extract environment and action information separately to train an action recognition model. Materzyńska et al. [11] proposed Something-Else, a compositional model that builds spatial-temporal interaction graphs between hands and objects, achieving strong performance on the Something-Something-v2 dataset. Although effective, these methods depend heavily on accurate hand and object detection, which can fail under occlusions or strong motion.
@@
RGB-based approaches [15, 1, 9, 5, 16, 19] rely solely on raw visual appearance for action recognition, without incorporating explicit motion or geometric cues. Many of these high-performing models are trained on large-scale datasets, often leveraging extensive pretraining and carefully designed architectures. In contrast, we deliberately adopt a simple model design to isolate and clearly demonstrate the effectiveness of incorporating 2D point tracks. Unlike RGB-based features that primarily encode texture and color statistics, point trajectories provide explicit geometric and temporal information, capturing object and camera motion directly from the image sequence. Girdhar et al. [5] propose the Anticipative Video Transformer, which models temporal dependencies in egocentric videos to predict upcoming actions. Their model implicitly attends to image regions containing hands and objects, and its performance improves with increased temporal context. However, the paper does not clarify whether motion is inferred implicitly through appearance changes between frames. In contrast, our approach introduces motion information explicitly by integrating 2D point tracks, enabling the model to reason directly about the displacement and dynamics of scene elements rather than relying on learned appearance differences. This explicit representation of motion complements the visual features and enhances interpretability in modeling fine-grained hand-object interactions and providing robustness to camera motion. Tong et al. [16] employ masked autoencoder pretraining to improve video representation learning, achieving strong performance on egocentric and other video benchmarks. Wang et al. [19] further extend pretrained video backbone models through an interactive prototype learning framework, enhancing active-object representations by leveraging actor motion cues within RGB inputs.
@@
Point tracking aims to estimate the trajectories of 2D points across video frames, providing fine-grained motion information that supports higher-level tasks such as action recognition and object interaction analysis. Recent advances in this area have focused on improving both tracking accuracy and computational efficiency. CoTracker [8] introduced a transformer-based architecture capable of jointly tracking multiple points across frames. Subsequent versions have further improved robustness, accuracy, and inference speed, establishing CoTracker as a state-of-the-art solution for general-purpose point tracking. In our work, we adopt CoTracker due to the public availability of pretrained weights, which allows for efficient and reliable integration of high-quality point trajectories into our framework. EgoPoints [4] proposed a dedicated benchmark for dense point tracking in egocentric videos. They specifically addresse the challenges of egocentric settings, such as strong camera motion and frequent occlusions, and provide a comprehensive evaluation protocol for tracking performance. In addition, the authors improve upon CoTracker by adapting it to egocentric motion patterns and by providing a large-scale dataset for benchmarking.
@@ 3 Action Recognition using Point Tracks
We present TRec, a transformer-based model for hand-object interaction recognition that jointly leverages 2D point tracks and image representations. TRec tracks both background and foreground points throughout a video to capture rich motion cues without relying on explicit object or hand detections. The model employs a lightweight image encoder to extract spatial features from individual frames and a transformer architecture [17] to integrate temporal dependencies across frames. A set of randomly sampled 2D points is initialized within each video, covering both foreground and background regions, and their trajectories are tracked over time. The transformer receives as input the encoded image features together with these 2D point tracks, enabling it to reason jointly about appearance and motion information. An additional multi-headed attention layer is used to aggregate the transformer outputs, and the resulting representation is passed through an multi-layer perceptron (MLP) based action recognition head to predict the action depicted in the video. We evaluate TRec on the Something-Something-v2 [6] dataset.
@@ 3.1 Dataset Curation and Preprocessing
Something-Something-v2. The Something-Something-v2 dataset [6] contains short video clips depicting human-object interactions recorded in a wide variety of everyday environments. It comprises 174 action templates describing generic activities of someone doing ”something”, such as pushing, pulling, or placing objects. It offers greater diversity in both visual context and action types, as well as a substantially large number of annotated samples, making it a valuable benchmark for training data-driven models. The dataset contains a diverse collection of videos, including egocentric recordings with significant camera motion as well as scenes captured from mostly stationary viewpoints. We use the dataset in its original form without any additional preprocessing or modifications, applying only standard data augmentation techniques.
Point Track Generation. We curate 2D point trajectories using CoTracker3 [7]. Figure 1 illustrates several examples of the generated point tracks. For each action sequence, we randomly sample 900 points covering the entire image, including both foreground and background regions. Video frames are sampled at 30 fps to obtain smooth trajectories. The resulting point coordinates are normalized following the procedure described in [2]. For each video we track a set of 900 points $p^{i}=\{(x^{i}_{t},y^{i}_{t})\in\mathbb{R}^{2}|t=0,...T\}$. Here, $i\in[0,P]$ denotes the $i^{\text{th}}$ point with $P$ being the total number of points while $t\in[0,T]$ refrs to the frame index, with $T$ denoting the number of frames in the video. We refer to a point at time step $t$ as $p_{t}^{i}\in\mathbb{R}^{2}$. The set of all point tracks within a video is denoted as $\mathbf{P}$.
(a) Throwing CD cover against sofa
(b) Put coaster into coaster holder
Figure 1: Trajectory of 2D points in example videos of the Something-Something dataset
@@ 3.2 Network Architecture
3.3 Action Recognition with Transformer
The transformer model receives as input the image features extracted by the encoder together with the corresponding 2D point tracks. This joint representation allows the model to integrate appearance and motion information in a unified latent space. Each frame contributes both spatial features from the image encoder and geometric motion cues from the tracked points. By processing these modalities jointly, the transformer learns to associate appearance changes with underlying motion dynamics, enabling robust recognition of fine-grained actions.
@@ 3.3 Action Recognition with Transformer
Finally, the pooled class token is passed through a lightweight MLP classification head to produce the final action prediction. During training, the model is optimized using a standard cross-entropy loss over the action classes. This simple yet effective design allows TRec to capture meaningful motion representations directly from tracked points and image features, achieving competitive performance while maintaining low computational complexity.
@@ 4.1 Track vs No Track
Table 1: Performance comparison with and without point tracks
@@ 4.3 Background Motion as Context for Action Recognition
Table 3: Model evaluation with KDE
@@ 5 Discussion
The experimental results demonstrate that incorporating 2D point tracks provides a clear advantage for egocentric action recognition compared to purely RGB-based methods. By explicitly encoding motion cues through tracked points, our model captures spatial-temporal dynamics that RGB alone cannot represent. This finding supports our initial hypothesis that motion trajectories offer a lightweight and interpretable alternative to computationally demanding approaches such as optical flow or hand-object detection.